# Performance and Storage Analysis of CRYSTALS-Kyber (ML-KEM) as a Post-Quantum Replacement for RSA and ECC

### v4

@@ Abstract
The steady advancement in quantum computer error correction technology has pushed the current record to 48 stable logical qubits, bringing us closer to machines capable of running Shor’s algorithm at scales that threaten RSA and ECC cryptography. While the timeline for developing such quantum computers remains uncertain, the cryptographic community must prepare for the transition to quantum-resistant algorithms. CRYSTALS-Kyber, standardized by NIST in 2022, represents a leading post-quantum cryptographic solution, but widespread adoption faces significant challenges. If this migration follows patterns similar to the SHA-1 to SHA-2 transition, organizations may experience prolonged periods of vulnerability, with substantial security and economic consequences. This study evaluates Kyber’s practical viability through performance testing across various implementation schemes, utilizing only standard built-in processor acceleration features (AES-NI, ASIMD) without any specialized hardware additions. Our findings demonstrate that Kyber provides robust security guarantees against quantum attacks while maintaining acceptable performance profiles for most contemporary applications, utilizing only commodity hardware with manufacturer-provided acceleration capabilities.
@@ 1 Introduction
The theoretical foundations established by Shor and Grover have evolved from academic concepts to practical concerns as quantum computing hardware continues to advance. While experts debate the timeline for achieving fault-tolerant quantum computers capable of running Shor’s algorithm at scale, the cryptographic community faces an urgent imperative: the transition to quantum-resistant algorithms cannot wait for quantum computers to become operational. Historical precedents in cryptographic transitions offer sobering lessons about the challenges ahead. The migration from SHA-1 to SHA-2, initiated in 2005 following the discovery of collision vulnerabilities, took over a decade to complete [PS], with many organizations maintaining vulnerable systems well beyond recommended timelines [PS]. This prolonged transition period exposed numerous systems to security risks and highlighted the substantial economic and operational costs associated with delayed cryptographic upgrades. Suppose the transition to post-quantum cryptography follows similar patterns. In that case, organizations may face extended periods of vulnerability to quantum attacks, with potentially catastrophic consequences for digital infrastructure, financial systems, and national security. In response to this challenge, the National Institute of Standards and Technology (NIST) initiated a comprehensive standardization process for post-quantum cryptographic algorithms, culminating in the selection of CRYSTALS-Kyber as the primary Key Encapsulation Mechanism (KEM) standard in 2022. Unlike RSA and ECC, which derive their security from number-theoretic problems vulnerable to quantum attacks, Kyber’s security relies on the learning with errors (LWE) problem over module lattices—a mathematical foundation believed to be resistant to both classical and quantum computational attacks. However, theoretical quantum resistance alone does not guarantee practical adoption. The success of any cryptographic standard depends critically on its performance characteristics, storage requirements, and compatibility with existing hardware infrastructure. Previous post-quantum proposals have faced significant barriers to adoption due to excessive computational overhead, prohibitive key sizes, or requirements for specialized hardware acceleration. To address these concerns and evaluate Kyber’s practical viability as a replacement for current cryptographic standards, this study conducts a comprehensive performance analysis across multiple architectures and implementation scenarios. This research contributes to the post-quantum cryptography transition by providing empirical evidence of Kyber’s performance characteristics under realistic deployment conditions. Our evaluation methodology utilizes standard hardware acceleration features available in commodity processors (AES-NI, AVX2, ASIMD) without requiring specialized, quantum-resistant hardware additions, ensuring that our findings accurately reflect the performance organizations can expect during real-world deployment. By comparing Kyber’s performance against equivalent-security implementations of RSA-7680 and SECP384R1 (ECC) across both x86_64 and ARM64 architectures, we establish benchmarks that inform migration planning and risk assessment for organizations preparing for the post-quantum era.
@@ 3 Cryptographic Schemes
To ensure a fair comparison, the following widely known and commonly used algorithms will be employed: RSA-7680, ECDH-SECP384R1 (ECC), and ML-KEM768 (Kyber). They all have the same security strengths according to the NIST [Bar06].
@@
To ensure consistency, the same DEM was used for all three schemes: the ChaCha20-Poly1305 authenticated encryption with associated data (AEAD) cipher [NL18]
@@ 5.1.2 Outgoing Shared Secret Derivation
It is important to note that for RSA, this “public-key” operation is significantly less costly than the “private-key” operation shown in the “Incoming Derivation” chart, which is an expected characteristic of the algorithm. Nevertheless, Kyber’s performance remains superior. For SECP384R1, the computational workload is identical for both the incoming and outgoing phases, which explains its consistently high cycle count. Kyber requires fewer than 300,000 cycles for this step, which reduces the latency of cryptographic handshakes relative to both RSA and ECC.

### v5

@@ Abstract
The steady advancement in quantum computer error correction technology has pushed the current record to 48 stable logical qubits, bringinga us a little closer to machines capable of running Shor’s algorithm at scales that threaten RSA and ECC cryptography. While the timeline for developing such quantum computers remains uncertain, the cryptographic community must prepare for the transition to quantum-resistant algorithms. CRYSTALS-Kyber, standardized by NIST in 2022, represents a leading post-quantum cryptographic solution, but widespread adoption faces significant challenges. If this migration follows patterns similar to the SHA-1 to SHA-2 transition, organizations may experience prolonged periods of vulnerability, with substantial security and economic consequences. This study evaluates Kyber’s practical viability through performance testing across various implementation schemes, utilizing only standard built-in processor acceleration features (AES-NI, ASIMD) without any specialized hardware additions. Our findings demonstrate that Kyber provides robust security guarantees against quantum attacks while maintaining acceptable performance profiles for most contemporary applications, utilizing only commodity hardware with manufacturer-provided acceleration capabilities.
@@ 1 Introduction
The theoretical foundations established by Shor and Grover [Sho94, Gro96] have evolved from academic concepts to practical concerns as quantum computing hardware continues to advance. While experts debate the timeline for achieving fault-tolerant quantum computers capable of running Shor’s algorithm at scale, the cryptographic community faces an urgent imperative: the transition to quantum-resistant algorithms cannot wait for quantum computers to become operational [Glo24, Gid25]. Historical precedents in cryptographic transitions offer sobering lessons about the challenges ahead. The migration from SHA-1 to SHA-2, initiated in 2005 following the discovery of collision vulnerabilities [WYY05, SBK+17, Nat15], took over a decade to complete [PS], with many organizations maintaining vulnerable systems well beyond recommended timelines [PS]. This prolonged transition period exposed numerous systems to security risks and highlighted the substantial economic and operational costs associated with delayed cryptographic upgrades. Suppose the transition to post-quantum cryptography follows similar patterns. In that case, organizations may face extended periods of vulnerability to quantum attacks, with potentially catastrophic consequences for digital infrastructure, financial systems, and national security. In response to this challenge, the National Institute of Standards and Technology (NIST) initiated a comprehensive standardization process for post-quantum cryptographic algorithms, culminating in the selection of CRYSTALS-Kyber as the primary Key Encapsulation Mechanism (KEM) standard in 2022 [Nat24]. Unlike RSA and ECC, which derive their security from number-theoretic problems vulnerable to quantum attacks, Kyber’s security relies on the learning with errors (LWE) problem over module lattices—a mathematical foundation believed to be resistant to both classical and quantum computational attacks [BDK+18]. However, theoretical quantum resistance alone does not guarantee practical adoption. The success of any cryptographic standard depends critically on its performance characteristics, storage requirements, and compatibility with existing hardware infrastructure. Previous post-quantum proposals have faced significant barriers to adoption due to excessive computational overhead, prohibitive key sizes, or requirements for specialized hardware acceleration. To address these concerns and evaluate Kyber’s practical viability as a replacement for current cryptographic standards, this study conducts a comprehensive performance analysis across multiple architectures and implementation scenarios. This research contributes to the post-quantum cryptography transition by providing empirical evidence of Kyber’s performance characteristics under realistic deployment conditions. Our evaluation methodology utilizes standard hardware acceleration features available in commodity processors (AES-NI, AVX2, ASIMD) without requiring specialized, quantum-resistant hardware additions, ensuring that our findings accurately reflect the performance organizations can expect during real-world deployment. By comparing Kyber’s performance against equivalent-security implementations of RSA-7680 and SECP384R1 (ECC) across both x86_64 and ARM64 architectures, we establish benchmarks that inform migration planning and risk assessment for organizations preparing for the post-quantum era.
@@ 3 Cryptographic Schemes
To ensure a fair comparison, the following widely known and commonly used algorithms will be employed: RSA-7680, ECDH-SECP384R1 (ECC), and ML-KEM768 (Kyber). They all have the same security bit strengths according to the NIST [Bar06].
@@
To ensure consistency, the same DEM —ChaCha20-Poly1305 authenticated encryption with associated data (AEAD) cipher [NL18]— was used for all three schemes.
@@ 5.1.2 Outgoing Shared Secret Derivation
It is important to note that, in RSA, this “public-key” operation is significantly less costly than the “private-key” operation shown in the “Incoming Derivation” chart, which is an expected characteristic of the algorithm. Nevertheless, Kyber’s performance remains superior. For SECP384R1, the computational workload is identical for both the incoming and outgoing phases, which explains its consistently high cycle count. Kyber requires fewer than 300,000 cycles for this step, which reduces the latency of cryptographic handshakes relative to both RSA and ECC.